### v1

# T-FAKE: Synthesizing Thermal Images for Facial Landmarking

Abstract
Facial analysis is a key component in a wide range of applications such as security, autonomous driving, entertainment, and healthcare. Despite the availability of various facial RGB datasets, the thermal modality, which plays a crucial role in life sciences, medicine, and biometrics, has been largely overlooked. To address this gap, we introduce the T-FAKE dataset, a new large-scale synthetic thermal dataset with sparse and dense landmarks. To facilitate the creation of the dataset, we propose a novel RGB2Thermal loss function, which enables the transfer of thermal style to RGB faces. By utilizing the Wasserstein distance between thermal and RGB patches and the statistical analysis of clinical temperature distributions on faces, we ensure that the generated thermal images closely resemble real samples. Using RGB2Thermal style transfer based on our RGB2Thermal loss function, we create the T-FAKE dataset, a large-scale synthetic thermal dataset of faces. Leveraging our novel T-FAKE dataset, probabilistic landmark prediction, and label adaptation networks, we demonstrate significant improvements in landmark detection methods on thermal images across different landmark conventions. Our models show excellent performance with both sparse 70-point landmarks and dense 478-point landmark annotations. Our code and models are available at https://github.com/phflot/tfake.
1 Introduction
*Equal contribution.11footnotetext: Systems Neuroscience & Neurotechnology Unit, Faculty of Medicine, Saarland University & htw saar22footnotetext: Institute of Mathematics, Technische Universität Berlin33footnotetext: Max Planck Institute for Informatics, Saarland Informatics Campus
Figure 1: Predictions of 70 (middle) and 478 (bottom) point facial landmarks on the challenging CHARLOTTE dataset which are estimated accurately by training a model with our synthetic data.
Functional thermal imaging is an acquisition modality with growing importance in many areas from life sciences over medicine to biometrics. Due to the close relation of body temperature with physiological states, thermal imaging of faces emerged as an important tool for screening infectious diseases [45, 40, 4] and for the analysis of neurocognitive and affective states [21]. As a consequence, there exists a growing need for thermal images with annotated facial landmarks. However, in contrast to RGB images, annotated data for thermal images of faces is still rarely available. For example, the largest thermal datasets with facial landmark annotations contain only tens to hundreds of individuals [29, 12, 43, 3], while the largest datasets for face detection for RGB images reached the mark of a million faces (of more than 500k individuals) with the MegaFace benchmark already in 2016 [24]. Moreover, large synthetic RGB datasets, like FAKE [52] appeared to be valuable for landmark detection training.
Generally, the detection of facial landmarks is an important downstream task for facial analysis. These 2D coordinates of corresponding points across faces are part of methods, e.g., for alignment of faces to extract regions of interest [53], the animation of 3D avatars [35], motion magnification [14, 13] or deepfakes [46]. The most common convention for facial landmark annotations uses 68 points across the face and face boundary, see, e.g. [48]. Even though the 68-point convention remains challenging for thermal images [42, 37], the importance of dense landmarkers with more than 300 points has been demonstrated [52, 53, 35].
Faces in thermal images share structural similarities with RGB images. However, temporally varying environmental factors such as temperature and rain droplets, and physiological factors such as blood perfusion changes and vascular details lead to much higher variability in the appearance of faces within individuals [20]. Sweating, flushing, or outdoor rain can result in time-dependent texture changes. Features that are very stable in RGB images such as the contrast of the nose and the eyes can become inverted when the nose or the face is cooled down. Due to the structural similarities, landmarkers trained on RGB datasets were also used under specific thermal imaging conditions [12]. However, they show high failure rates and produce imprecise landmark predictions on challenging thermal images. Up to now, there exist only a few sparse thermal landmark detection methods [29, 37, 43, 33, 32], and we are unaware of a dense thermal or a multimodal RGB+Thermal landmark system. In this paper, we aim to address this gap by introducing the T-FAKE dataset, the first large-scale synthetic thermal dataset featuring both sparse and dense landmarks for facial analysis. We propose to thermalize a synthetic RGB dataset using our novel RGB2Thermal loss function, which comprises three key components: i) a supervised data term that controls the generation of thermal faces based on a small subset of RGB-thermal pairs, ii) a Wasserstein distance term that aligns the patch distributions of the generated synthetic thermal images with those of real thermal images, and iii) a term incorporating prior information on clinical temperature statistics for different facial regions. To validate the quality of our T-FAKE dataset, we employ a probabilistic landmark prediction method by minimizing the negative log-likelihood function of the landmarks which are assumed to follow a Gaussian distribution as proposed in [53]. Moreover, by introducing a learnable adapter between different landmark conventions, we unify various methods and datasets for evaluation.
We thoroughly evaluate the landmark prediction method trained on our T-FAKE dataset on the established RGB dataset 300W [48] and the challenging thermal dataset CHARLOTTE [3], which includes controlled modulation of environmental parameters such as pose and resolution. Our model demonstrates superior performance on the thermal landmark prediction benchmark compared to previous methods while maintaining performance on the RGB images comparable to state-of-the-art RGB methods. Fig. 1 illustrates the performance of our model for sparse and dense landmarking. In summary, we provide the following contributions:
  • The first large-scale facial synthetic thermal dataset T-FAKE with sparse and dense landmark annotations;
  • A novel RGB2Thermal loss to facilitate thermal image synthesis overcoming limited lab-recorded training data;
  • Training of the up-to-date first dense thermal landmarker and a state-of-the-art multimodal RGB+Thermal sparse landmarker in combination with highly structured benchmarking across different landmark conventions and modalities.
2 Related Work
[…]
Aachen[29] 2018 68 2935/90 90
ARL-MMFD [54] 2019 (request) 6 ?/111 250, 500, 750
DRIVE-IN [12] 2021 68 and 468 ?/436 $\sim$80-150
ARL-VTF [43] 2021 68 ?/395 210
TFW [32] 2022 5 9982/147 100-?
SF-TL54 [33] 2022 54 2556/142 100
CHARLOTTE [3] 2022 73 (43 profile) 7684/10 100-660
T-FAKE (ours) 2024 70 and 468 200k/2k $\sim$80-150
Table 1: Comparison of thermal landmark datasets. Number of frames (F) and number of subjects (S) are indicated as well as distance to the camera if known. Unavailable data are marked by a question mark.
Facial RGB landmarking. The most widely used annotation convention defines 68-point landmarks which are available for many facial alignment datasets [27, 55, 6]. Sometimes points on the eye or other details are additionally added [52, 3]. While an in-depth review of facial RGB landmarking is beyond the scope of this work, we introduce those that are relevant to thermal landmarking methods. Kazemi et al. [23] introduced a traditional computer vision method for landmark prediction that is still widely used via the dlib library [25]. Kowalski et al. [30] proposed the deep alignment network (DAN) which uses convolutional neural networks and iteratively refines landmark predictions. Around the same time, Bulat et al. [6] proposed the facial alignment network (FA). Both methods predict heat maps of landmark positions from which coordinates are computed. Explicit integration of uncertainty quantification into a deep landmark prediction model via the LUVLi loss has been proposed by Kumar et al. [31]. In recent years, the quality of image annotation became a limiting factor of modern algorithms which led to the emergence of synthetic data with high-quality ground truth [39]. Wood et al. [52, 53] trained landmarkers on synthetic data that outperform traditional approaches and proposed the FAKE dataset with 100k frames of 2k individual faces.
Thermal landmarking. Many tailored approaches for thermal landmarking use similar architectures as RGB landmarkers and are mainly developed alongside the publication of new datasets, see Table 1. Kuzdeuov et al. [32] released the TFW dataset with a total number of 9982 images from 147 subjects annotated with 5 facial landmarks and bounding boxes and trained YOLOv5 networks for landmark prediction. The SF-TL54 dataset [33] consists of 54 landmarks and 2556 frames of 142 subjects and has been used for landmark prediction with a dlib and a U-net model. Kopaczka et al. [29] also use a dlib model for their AACHEN dataset with annotated 2935 images of 90 subjects and 68-point landmarks. A highly structured dataset is the CHARLOTTE thermal face dataset with 10 subjects and 10k images by Ashrafi et al. [3], who annotated their dataset with 72-point landmarks. In contrast to these publicly available datasets, datasets with limited or no public availability include the the ARL-VTF dataset [43] (500k frames of 395 subjects) and a dataset recorded in SARS-CoV-2 drive-in stations (DRIVE-IN) [12] (six minutes recordings at $50$Hz of 436 subjects). The authors of the DRIVE-IN dataset found that a stack of pre-processing functions already leads to a good performance of existing RGB landmark models on thermal images. Similarly, Poster et al. [44] proposed to use transfer learning on a siamese DAN to improve thermal landmark prediction with RGB facial landmarks. Mallat et al. [37] used RGB2Thermal image translation (‘thermalization’), active appearance models, and deep alignment networks of RGB datasets for thermal landmark prediction. In contrast to the previous work, our dataset contains thermal images of 2k individuals that are annotated with sparse 70 as well as dense 478 point landmarks.
Thermalization. In addition to thermal landmark datasets, various datasets of paired and aligned thermal and RGB images without annotations exist. These can be employed for the training of thermalization models. Mallat et al. thermalized facial images by minimizing a perceptual loss [37] using the VIS-TH dataset [36]. Similarly, generative adversarial networks (GANs) have been used for thermalization [54, 20] using the NCKU-VTF [20] dataset and the ARL-MMDF [54] dataset containing facial image data. In contrast to the limited public availability of these two datasets, a publicly available dataset with paired facial images is the multi-modal disguise Sejong face database (SEJONG) with 100 subjects, 15 disguise variations and 4849 RGB+Thermal image pairs by Cheema and Moon [7]. Thermalization of general scenes has been achieved by Kniaz et al. [26] with a modified Pix2Pix GAN, where they used the publicly available ThermalWorld dataset with 5098 image pairs displaying various object classes. In contrast to previous works, we employ a semi-supervised RGB2Thermal loss that incorporates a semantic segmentation to improve generalization.
Figure 2: Three steps of our training pipeline (models and their loss functions): 1. thermalization network $\mathcal{T}_{\theta}$ with loss (1) - (3). 2. landmark prediction network $\mathcal{T}_{\psi}$ with (6). 3. landmark adaptation network $\mathcal{T}_{\zeta}$ with (7). We show black dots for thermal landmarks and white dots for RGB landmarks. Note the landmark adaptation on the forehead in the last step.
3 T-FAKE Dataset
Recently, the success of landmarkers trained solely on the synthetic facial landmarking dataset FAKE [52] has demonstrated the potential of synthetic data for landmarking. We aim to transfer this success story to the largely overlooked thermal domain and the multimodal setting by introducing our synthetic thermal landmarking dataset T-FAKE featuring 2k unique subjects.
In particular, we train a semi-supervised image translation model for the thermalization of RGB images using real and synthetic data. To account for variations in the data in terms of poses, hairstyles, and expressions, we employ a semi-supervised loss.
3.1 Thermalization
To generate our large-scale synthetic T-FAKE dataset, we propose a thermalization framework. First, we train our network in a semi-supervised manner based on training pairs of RGB+Thermal images $(\mathbf{X}_{\text{ RGB}},\mathbf{X}_{\text{T}})=(X_{\text{ RGB}}^{n},X_{\text{T}}^{n})_{n=1}^{N}$ from the lab-recorded SEJONG dataset. Further, we use the synthetic RGB images from FAKE $\mathbf{X}_{\text{FAKE}}=(X_{\text{FAKE}}^{m})_{m=1}^{M}$ for which no thermal counterpart exits. Indeed the thermalization of these data appears to be difficult due to the large differences between real-world lab-recorded and synthetic in-the-wild data when facing thermalization tasks. For RGB2Thermal translation, we optimize a U-net [47] $T_{\theta}$ by minimizing the loss function
$\mathcal{L}(\theta)$ $:=\frac{1}{N}\sum_{n=1}^{N}{\|T_{\theta}(X_{\text{ RGB}}^{n})-X_{\text{T}}^{n}\|}_{2}^{2}$ (1)
$+\lambda_{W}\operatorname{W}(T_{\theta}(\mathbf{X}_{\text{FAKE}}),\mathbf{X}_{\text{T}})$ (2)
$+\lambda_{R}\frac{1}{M}\sum_{m=1}^{M}\operatorname{R}(T_{\theta}(X_{\text{FAKE}}^{m}),\quad\lambda_{W},\lambda_{R}>0.$ (3)
This loss allows us to simultaneously train a domain-adaptive image translation model from paired and unpaired data. The first summand (1) penalizes the deviation of the generated thermal image from the true one and serves as a reconstruction loss for known image pairs. The second summand (2) exploits the Wasserstein distance on the patch distribution of thermalized FAKE images and natural thermal images. This style loss ensures that a generated image displays the thermal image statistics. To be more precise, recall that the squared Wasserstein-2 distance $\mathcal{W}_{2}^{2}$ between empirical measures $\mu=\frac{1}{K}\sum_{k=1}^{K}\delta_{x_{k}}$ and $\nu=\frac{1}{L}\sum_{l=1}^{L}\delta_{y_{l}}$ is given by
$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\pi\in\Pi}F(\pi):=\sum_{k=1}^{K}\sum_{l=1}^{L}\|x_{k}-y_{l}\|^{2}\pi_{k,l},$ (4)
where $\Pi$ denotes the set of non-negative $K\times L$ matrices which rows and columns sum up to $1/K$ and $1/L$, respectively. Calculation with this distance requires solving costly optimization problems. As a remedy, the entropy-regularized Wasserstein distance
$\mathcal{W}_{2,E}^{2}(\mu,\nu)$ $:=\inf_{\pi\in\Pi}\big{\{}F(\pi)+\lambda_{E}\sum_{k=1}^{K}\sum_{l=1}^{L}\pi_{k,l}\log\left(\pi_{k,l}\right)\big{\}}$
with some (small) regularization $\lambda_{E}>0$ has been proposed [9]. The new formulation and the resulting gradients can efficiently be calculated via the Sinkhorn algorithm [11, 41]. To use this loss for thermal translation, we focus on the patch distributions of the generated and real images. For that purpose, we now extract (overlapping) image patches $P_{i}X:=X[i_{1}:i_{1}+s,i_{2}:i_{2}+s]$, $i=(i_{1},i_{2})\in I$ of size $s\times s$ and consider the empirical measures generated by the patch distributions of $T_{\theta}(\mathbf{X}_{\text{FAKE}})$,
$\mu_{\text{FAKE}}:=\tfrac{1}{N}\tfrac{1}{|I|}\sum_{n=1}^{N}\sum_{i=1}^{|I|}\delta_{P_{i}\left(T_{\theta}(X_{\text{FAKE}}^{n})\right)},$
and similarly $\mu_{\text{T}}$ from $X_{\text{T}}$. Now we penalize in (2) the term
$\operatorname{W}(T_{\theta}(\mathbf{X}_{\text{FAKE}}),\mathbf{X}_{\text{T}})$ $:=\mathcal{W}_{2,E}^{2}(\mu_{\text{FAKE}},\mu_{\text{T}}).$ (5)
During training, we apply a multiscale approach by applying this regularizer over multiple image scales and summing over the results. As a result, the patch statistics of downscaled images are incorporated into the empirical loss. The proposed patch distribution matching-based regularization was inspired by its application in related tasks such as image reconstruction [1, 41], style transfer [34], texture reconstruction [18, 2], and image translation [22].
Finally, the distribution of temperature across different regions of the face and body are well-studied [50, 8], e.g., warmer areas around the eyes, colder scalp hair and glasses blocking thermal heat. Thus, the third summand (3) incorporates such knowledge about the expected temperature in different facial areas. To be specific, we make use of a given semantic segmentation of the synthetic dataset of the image $\mathbf{X}_{\text{F-RGB}}$ into pixels corresponding to 18 regions $S_{i}$, e.g., ‘background’, ‘nose’, ‘headwear’ and so on. Moreover, let $T_{i}$ be the average temperature of the hardcoded reference temperatures in the region $S_{i}$ based on measurements presented in empirical studies [8, 50, 3]. Then, we compute
$\operatorname{R}\left(T_{\theta}(\mathbf{X}_{\text{FAKE}})\right)=\sum_{m=1}^{M}\sum_{i=1}^{18}\omega_{i}(\overline{S_{i}(T_{\theta}(X_{\text{FAKE}}^{m})})-T_{i})^{2},$
where the weights $\omega_{i}$ are proportional to the region sizes and $\overline{S_{i}}$ is the mean temperature in the $i$-th region. Note that facial temperatures depend on the environmental conditions. Hence, we train $T_{\theta}$ two times with two different reference temperature sets to simulate ‘cold’ and ‘warm’ environmental conditions. The different regularizers are visualized in Fig. 2 (left) and the segmentation masks are shown in Fig. 3 (bottom).
Predictions of the normalized temperature with our model compared to a model without regularization are shown in Fig. 3. The advantages of our regularization in a domain-adapted setting are clearly visible. The use of regularization reduces image artifacts. The patch regularizer (2) promotes natural texture synthesis and because of our segmentation regularizer (3) temperature differences between facial areas are more pronounced. Hence, our method allows for the adaptation of images with shadows, various skin colors, and distinctive facial expressions. In particular, such features may be underrepresented in facial thermal imaging datasets recorded in lab conditions. Note that the resulting domain adaptation needs for synthetic data and the access to ground truth segmentations differentiates our task from the thermalization tasks in [26, 54, 37, 20].
Figure 3: Frames from the FAKE dataset [52] before (first row) and after thermalization without any regularization (second row) and our domain-adaptive model (third row). The last row displays the segmentation masks with gray levels which are proportional to the reference temperature values used in regularizer (3). The first three images exemplify the ‘cold’ setup with colder noses and the last two the ‘warm’ setup with warmer noses.
3.2 The Dataset
Having trained $T_{\theta}$, we use it to thermalize the FAKE dataset which results in the thermal dataset T-FAKE. For each thermal condition, the dataset contains 100k images featuring 2k unique subjects. Each image is available in a simulated ‘cold’ and a ‘warm’ environment adding up to 200k images. This number of frames and subjects surpasses the numbers in other thermal landmarking datasets [29, 54, 12, 43, 32, 33, 3] known to the authors. In contrast to thermal landmarking datasets with a fixed distance between the subjects and the camera [29, 43, 33], the distance varies within an estimated range of 80cm to 150cm for the T-FAKE dataset. Moreover, the synthesized images are based on facial scans with a wide range of age and ethnicity and a balanced number of people identifying as male or female [52]. In addition to the original ground truth sparse landmark annotations, we add dense annotations by transferring Mediapipe [35] predictions for the FAKE to the T-FAKE dataset. This makes our dataset the first densely annotated thermal landmarking dataset. While we found many of the large thermal landmarking datasets to be non-public [54, 12, 43], our complete dataset with both thermal conditions and annotations is available for download. In combination with the original FAKE dataset, we can train multimodal facial landmarkers jointly on RGB and thermal images as described in the next section. Table 1 displays a detailed comparison between thermal landmarking datasets
4 Facial Landmarking
Learning with synthetic images has proven to be very useful in facial landmarking [54, 52, 53]. Based on the proposed T-FAKE and the original FAKE dataset, we can thus train a supervised thermal or multimodal landmark prediction model for thermal and RGB images using synthetic data only. Moreover, we can adapt our model to new annotation conventions by training a small-scale label adaptation model. The data flow for these two models is visualized in Fig. 2.
4.1 Landmark Prediction
[…]
$-\log p(y|\mu,\sigma)=\log\left(2\pi\sigma^{2}\right)+\frac{\|\mu-y\|_{2}^{2}}{2\sigma^{2}},$
the authors of [53] proposed to learn the network $T_{\psi}:=(T_{\psi}^{1},T_{\psi}^{2})$ with $X\mapsto(\boldsymbol{\mu},\boldsymbol{\sigma}^{2})$, where $\boldsymbol{\mu}=(\mu_{l})_{l=1}^{L}\in(\mathbb{R}^{2})^{L}$ and $\boldsymbol{\sigma}^{2}=(\sigma_{l}^{2})_{l=1}^{L}\in\mathbb{R}^{L}$ by minimizing the loss function
$\mathcal{L}(\psi)$ $:=\sum_{l=1}^{L}\Big{(}\sum_{j=1}^{J}\log T^{2}_{\psi,l}(X^{j})+\frac{\|T^{1}_{\psi,l}(X^{j})-y_{l}^{j}\|_{2}^{2}}{2\,T^{2}_{\psi,l}(X^{j})}\Big{)}.$ (6)
During training, we clip $T^{2}_{\psi,l}\geq\varepsilon>0$ to avoid numerical instability due to too small variances. The first sum controls the location accuracy and the second one the uncertainty prediction. Having learned the network $T_{\psi}$, it produces not only the expected values $\boldsymbol{\mu}$ of the corresponding landmarks, but also a measure for its uncertainty $\boldsymbol{\sigma}^{2}$.
4.2 Label Adaptation
A common approach that is especially crucial for synthetic trained data is label adaptation [52, 53]. It allows us to adapt to different landmark conventions between the training dataset and an evaluation dataset. Given two sequences $(\hat{\mathbf{y}}_{k},\mathbf{y}_{k})_{k=1}^{K}$ with predicted landmarks $\hat{\mathbf{y}}_{k}\coloneqq(\hat{y}^{l}_{k})_{l=1}^{\hat{L}}\in\mathbb{R}^{2\hat{L}}$, and with ground truth landmarks $\mathbf{y}_{k}:=(y^{l}_{k})_{l=1}^{L}\in\mathbb{R}^{2\hat{L}}$ with possibly $\hat{L}\neq L$, we aim to train a third label adaptation model $T_{\zeta}:\mathbb{R}^{2\hat{L}}\to\mathbb{R}^{2{L}}$. Here, we use the loss
$\mathcal{L}(\zeta)=\frac{1}{K}\sum_{k=1}^{K}\|T_{\zeta}(\hat{\mathbf{y}}_{k})-\mathbf{y}_{k}\|_{1}.$ (7)
Since we deal with low-dimensional data, we use a simple multi-layer fully connected perceptron with 5 layers for this task, see supplementary material for details.
5 Experiments
[…]
5.1 Implementation Details
Thermalization. We train on the RGB+Thermal SEJONG dataset recorded in lab conditions and the large-scale synthetic FAKE dataset featuring in-the-wild RGB images. We train on a resolution of $256\times 256$. For the patch-based regularizers, we use a patch size of $8$. We disregard the background by excluding synthetic patches based on the ground truth background segmentation and completely black real patches. For our multi-scale approach, we sum the regularizer over $5$ scales with a downsampling factor of $0.5$. Furthermore, facial temperatures depend on the surrounding temperature.
Thus, we train a ‘warm’ and a ‘cold’ model with different reference temperature values for the segmentation-based regularizers, see Fig. 3. For data augmentation, we use the same random rotations and cropping for the natural RGB and thermal images. Moreover, we apply random color changes, blurring, and shadow augmentations [38] exclusively to the natural RGB images. Next, we apply random rotations and cropping for the synthetic RGB images. Here, we also fill holes in the original ‘glasses’ segmentation masks showing outlines of the frame only to highlight transparent, but heat-blocking glass or plastic, see Fig. 3 (bottom right). Lastly, we replace the original background with a black background based on the known semantic segmentation. We apply our two models to the FAKE dataset at the original $512\times 512$ resolution to generate the T-FAKE dataset. This results in the up-to-date largest thermal landmarking dataset with 200k images of 2k unique subjects and two thermal settings, see Table 1.
Landmarker. We simultaneously train our landmark prediction model using a sparse 70-point and dense 468-point landmark convention with an image size of $224\times 224$ on our new T-FAKE thermal dataset and the original FAKE RGB dataset. Additionally, we include random landmark positions on a texture dataset [16] as negative examples into the dataset to increase the learned uncertainty $\sigma^{2}$ on images without faces. Due to the limited, public availability of the dense ground truth in the FAKE dataset [52], we use the 468-point predictions of the MediaPipe model for the synthetic RGB images as our dense ground truth dense annotations. We make use of spatial augmentations, including random shear, rotations, resizing, and cropping to allow the landmarker to learn a large variation of face orientations without the need for a dedicated face detector. Furthermore, we employ Gaussian blur, color jitter, and global pixel offset for thermal images during training. On thermal images, we fill in the background with random textures from the texture dataset [16]. During inference, we use a multi-scale sliding window evaluation to generalize our model to varying image sizes. We downsample iteratively with a factor of $0.75$ until the height or the width is smaller than $224$. For each image scale, we run our model on sliding windows of size $224\times 224$ with a stride of $20$. For our final landmark prediction, we pool all predictions and use the landmark with the smallest predicted standard deviation across all scales and all sliding windows.
Label Adaptation. Our label adaptation network is trained solely on the two-dimensional landmark predictions and ground truth annotations of a training split of a benchmark dataset and not on the images. This dataset is later excluded from the final evaluation. We do not use the predicted standard deviation at this stage. Again, we use random rotations and shear transformations during training.
Evaluation Metrics for Landmarking. We report the normalized mean error (NME) [52, 48, 56] for evaluation, i.e., the mean absolute error (MAE) normalized with an image-dependent constant to account for varying image sizes. Often, normalization is performed by the distance between the outermost eye landmarks (NME Interocular) [52, 48]. However, the CHARLOTTE benchmark dataset contains many examples with side profile views where only one of the eyes is visible. Therefore, we use the average height and width of the bounding box around the face for normalization [56] (NME W/H). Apart from the NME, we report the failure rates for each of the reference methods. The failure rate is the percentage of frames where no faces could be found. Frames without recognized faces are not included nor weighted in the NMEs. The method of rejection varies for each landmarker. We use vanilla parameters except for MediaPipe [35], where we set the minimal detection confidence to $0.01$. For our method, we evaluate it i) on all frames without failure rate and ii) on all frames where the average uncertainty scores $\sigma_{l}$ of the predicted landmarks $l$, $1\leq l\leq L$, is less than $\bar{\sigma}=6\times 10^{-4}$.
Metric Method Charlotte High Charlotte Low Charlotte Side Charlotte Front Charlotte Full
NME Aachen [28] 0.1677 0.2405 0.2942 0.1731 0.2054
3FabRec [5] 0.2862 0.3195 0.1810 0.4259 0.3033
3FabRec* 0.2646 0.3272 0.1921 0.4015 0.2967
Mediapipe [17] 0.1900 0.2745 0.2879 0.1865 0.2262
Mediapipe* 0.1400 0.2621 0.1838 0.2164 0.2023
FA [6] 0.1070 0.2790 0.1222 0.2508 0.1875
FA* 0.1007 0.2207 0.1075 0.2123 0.1593
TFW [32] 0.0787 0.1249 0.0701 0.1311 0.1012
SF-TL54 U-Net [33] 0.2476 0.3134 0.1757 0.3596 0.2829
SF-TL54 Dlib [33] 0.2159 0.2174 0.1365 0.2741 0.2167
Ours dense ($\bar{\sigma}<\infty$) 0.0805 0.1285 0.0846 0.1258 0.1050
Ours ($\bar{\sigma}<\infty$) 0.0740 0.1346 0.0684 0.1420 0.1051
Ours ($\bar{\sigma}<6\times 10^{-4}$) 0.0729 0.1139 0.0628 0.1219 0.0923
Failure rate (%) Aachen [29] 49.32 48.72 72.85 25.15 49.01
Mediapipe [35] 25.05 46.85 50.09 22.40 36.25
Mediapipe* 12.80 14.07 24.97 1.92 13.45
FA [6] 12.95 27.43 21.71 19.07 20.39
FA* 0.83 10.45 4.75 6.80 5.77
TFW [32] 13.14 21.55 19.26 15.66 17.46
SF-TL54 U-NET / Dlib [33] 29.29 22.57 38.16 13.50 25.84
Ours ($\bar{\sigma}<6\times 10^{-4}$) 1.63 16.45 9.15 9.34 9.25
Table 2: Results on the CHARLOTTE dataset. Pre-processing with the full pre-processing stack for RGB landmarkers is indicated by *. The low resolution (3434 frames) and high resolution (3250 frames) are defined by the median image width of 200 pixels. We denote the confidence threshold with $\bar{\sigma}$. Note that we include all predictions for $\bar{\sigma}<\infty$ and 3FabRec resulting in a failure rate of 0 by design.
Figure 4: Results on examples from the CHARLOTTE dataset [3] of FA and Mediapipe with the full pre-processing stack and our models without label adaptation. While the RGB methods can handle some challenging poses, they produce broken landmarks on certain difficult images.
5.2 Benchmark Datasets for Landmarking
To compare thermal performance, we use the CHARLOTTE dataset due to its high variability in resolution, pose, and background temperature. The dataset is annotated with either 72 landmarks for frontal images or 43 landmarks for lateral images. The 43-point landmarks represent the visible landmarks in profile pictures with our information on orientation. The dataset contains a total of 7685 annotated frames with different resolutions with a minimal one of around 70x60. For evaluation, we split the CHARLOTTE dataset based on image size and perspective. The CHARLOTTE High condition contains all images larger than the median image width of 200 pixels and CHARLOTTE Low all smaller images. The condition CHARLOTTE Side includes all images with profile annotations (43 landmarks), and CHARLOTTE Front contains all frontal faces (73 landmarks). In addition, we evaluate on the established RGB landmark dataset 300W [48] to verify RGB landmark prediction performance. It uses the classical 68-point convention. Here, we crop images to the pre-defined bounding boxes around each face for our evaluation. The average image resolution is about 290 pixels.
300W
Method NME W/H NME Interocular
3FabRec [5] 0.0376 0.0646
FA [6] 0.0618 0.1083
Ours 0.0330 0.0577
Ablation Study
GB 0.0303 0.0523
Table 3: Comparison of RGB performance of our method. We trained the label adaptation on 300 random samples of the combined W300 outdoor and indoor scenes (600 images) and evaluated the remaining images.
Regularization FID
$\lambda_{R}=0,\lambda_{W}=0$ 0.3325 $\pm$ 0.0039
$\lambda_{R}=0,\lambda_{W}=0.01C$ 0.3139 $\pm$ 0.0035
$\lambda_{R}=1,\lambda_{W}=0$ 0.2719 $\pm$ 0.0054
$\lambda_{R}=1,\lambda_{W}=0.01C$ 0.2666 $\pm$ 0.0037
Table 4: Influence of the regularizers on the thermalization. The parameter $\lambda_{W}$ corresponds to the impact of the patch-based regularizer (Eq. (2)) and $\lambda_{R}$ corresponds to the segmentation-based regularizer (Eq. (3)). We use $C=(5\cdot 8^{2})^{(-1)}$.
Metric Method Charlotte High Charlotte Low Charlotte Side Charlotte Front Charlotte Full
NME RGB 0.1055 0.2675 0.1241 0.2534 0.1887
RGB* 0.0933 0.2682 0.1312 0.2348 0.1824
Thermal Only 0.0832 0.1334 0.0677 0.1503 0.1090
Table 5: Ablation results on the CHARLOTTE dataset. Pre-processing with the full pre-processing stack for RGB landmarkers is indicated by *. Here, we do not exclude low-confidence images. Hence, we report the NME on all images without failure rate
5.3 Comparison to State-of-the-Art
We compare the performance of our landmarker trained on our full synthetic datasets against the state-of-the-art models on the CHARLOTTE dataset. Here, we include not only relevant thermal landmark detection systems but also RGB landmark detection systems. A previous study has shown that the performance of RGB landmarkers on thermal images can be significantly boosted by applying a stack of unlearned linear filters before prediction [12]. Consequently, the application of this pre-processing stack of linear filters allows us to include RGB landmarkers in our benchmark and to draw a fair comparison beyond task-specific thermal landmarkers. Because the CHARLOTTE dataset uses its own landmark convention, we use label adaptation for all methods for the image evaluation. We train each label adaptation model with a 1000-image training split on CHARLOTTE similar to [52] and exclude the training split from the final evaluation. Note that this approach allows us to include our sparse model and our dense model in the evaluation. To benchmark the RGB performance of the multimodal model, we evaluate our sparse model predictions against state-of-the-art RGB sparse landmarkers on the 300W dataset.
CHARLOTTE performance. The NME performance as well as the failure rates of the different methods varies drastically for the different CHARLOTTE splits. We observe the lowest performance across the landmarkers on the Front condition, followed by the low-resolution images. The FA and Mediapipe RGB methods show a performance that is comparable with the SF-TL54 and Aachen thermal landmarkers. Using a pre-processing stack [12], they consistently outperform them on the High condition in terms of NME and on Full, Front, Low and Side in terms of failure rate. Our model achieves state-of-the-art performance for sparse thermal landmarking across all different conditions. The only model with comparable performance is TFW which, however, only predicts a set of 5 landmarks. While the NME of TFW is slightly lower than our method for $\bar{\sigma}<\infty$ for Low, Front, and Full, our method outperforms TFW on all dataset splits in terms of NME and failure rate for $\bar{\sigma}<6\times 10^{-4}$. These results are presented in Table 2 and examples are visualized in Fig. 4.
300W performance. Additionally, we evaluate the RGB models on the 300W dataset in Table  3. Our model, trained on a combination of thermal and RGB facial images, demonstrates competitive performance on RGB images, being outperformed only by the original model trained exclusively on RGB data. The overall performances are consistent with the literature [52].
5.4 Thermalization Ablation Study
We ablate our different regularizers in Eq. (1)-(3) using the Fréchet Inception Distance (FID) [19]. The FID measures image distribution similarity between a real and a synthetic set of images based on an Inception v3 [51] model pre-trained on ImageNet [10]. For our evaluation, we report the average FID between the complete SEJONG dataset and twenty equal-sized samples from the FAKE dataset. Here, we use the SEJONG dataset as a reference thermal dataset to evaluate the similarity of differently thermalized images to true thermal images. We thermalize each sample with i) our model without regularization, ii) our model with patch-based regularization (Eq. 2) only, iii) our model with segmentation-based regularization (Eq. 3) only, and iv) our final model. All models are trained with the same setup and we randomly choose the ‘warm’ or the ‘cold’ variant for each image. Table 4 shows the average FIDs and the standard deviations. We see that the segmentation-based regularizer leads to a large boost in perceptual similarity and that the patch-based regularizer leads to a smaller boost. The combination of both regularizers gives the best result and confirms the effectiveness of our RGB2Thermal loss.
5.5 Landmarking Ablation Study
To study the impact of our thermal data, we additionally report the CHARLOTTE results of our landmarker trained with i) RGB images only and ii) thermal images only. Again, we use label adaptation for all variations. Table 5 shows the results. Training with the T-FAKE dataset significantly improves the accuracy of thermal landmarking across all conditions. In addition, multimodal training with the FAKE and T-FAKE datasets leads to better thermal landmarking performance than training only with the T-FAKE dataset. Moreover, we ablate our RGB performance on 300W by testing the model trained on RGB images only, see Table 3. Here, we see a slight drop in performance caused by multimodal training. Nevertheless, our multimodal model still outperforms FA [6] and 3FabRec [5].
6 Conclusion
To overcome data limitations in thermal landmarking, we established a semi-supervised RGB2Thermal image translation method. Using patch-based and segmentation-based regularization, we ensured the generalization of our thermalizer to the synthetic FAKE dataset. We employed this model to generate the up-to-date largest thermal landmarking dataset, the T-FAKE dataset. By training a probabilistic landmarker with the T-FAKE dataset, we established the up-to-date first dense thermal landmarker and a state-of-the-art RGB+Thermal sparse landmarker. We tested these landmarkers with strict benchmarking by including established thermal and RGB methods and using label adaptation for each benchmarked method. Our work shows that synthetic data is sufficient for the training of thermal landmarkers and can be done on multiple imaging modalities.
7 Acknowledgements
This study has partially been funded by the Federal Ministry of Education and Research (BMBF, grant numbers 13N15753 and 13N15754). M.P. acknowledges funding from the German Research Foundation (DFG) within the GRK2260 BIOQIC project 289347353. The authors acknowledge HPC resources support with hardware funded by the DFG within project 469073465. The authors would like to thank Daniel J. Strauss for discussions and Mayur Bhamborae and Pascal Hirsch for technical support.
Appendix A Thermalization
A.1 Reference Temperature Values
As described in the main document, we train two versions of the thermalizer $T_{\theta}$ to model facial temperature variations under different environmental conditions. For that purpose, we use two sets of reference temperatures, a ‘cold’ and a ‘warm’ condition, for the different facial regions that guide the segmentation-based regularizer. These are presented in Table 6. Thermal facial contrast is increased and overall body temperature is decreased for the ‘cold’ condition in comparison to the ‘warm’ condition which is in line with empirical findings [3].
Segmentation ‘Cold’ ’Warm’
Background, Glasses $<$20 $<$20
Skin 33 35
Nose 31.5 35
Eyes 34 35
Brows 31 34
Ears 32 35
Mouth Interior 35 35
Lips 32.5 35
Neck 34 35
Hair 30 30
Beard 31 32
Clothing 30 32
Headwear, Facewear 28 28
[…]
A.2 Thermalization Implementation Details
We use a U-Net $T_{\theta}$ with a Resnet34 encoder pre-trained on ImageNet [10] and train it for 10 epochs with SEJONG batches of size $64$ and FAKE batches of size $64$. Further, we use an Adam optimizer with an initial learning rate of 0.001 which we reduce to 0.0001 after 4 epochs. Based on a random split, we use 80% of the SEJONG data for training and all available FAKE data. Also, we use the geomloss [11] implementation for the Wasserstein patch loss with $\lambda_{E}=0.1^{6}$. We normalize the squared error loss $\|\cdot-\cdot\|_{2}^{2}$ by dividing with the image dimension, here $256^{2}$. Based on the $5$ scales and patch dimension $8^{2}$, we set $\lambda_{W}=0.01C$ for our final model with the normalization constant $C=(5\cdot 8^{2})^{-1}$. We set $\lambda_{R}=1$. Note that the choice of $\lambda_{R}=1$ and $C$ are motivated to control the value range. The (normalized) MSE data fidelity term $\|\cdot-\cdot\|_{2}^{2}$ which we evaluate on normalized SEJONG images takes values in the range $[0,1]$ for arbitrary images with pixel range $[0,1]$ due to dimensional normalization. To have a similar value range for the evaluated FAKE images, the patch-based regularizer $W$ takes values in $[0,1]$ for $\lambda_{W}=C$ with arbitrary normalized images. Moreover, the segmentation-based regularizer $R$ also takes values in $[0,1]$ for such images given arbitrary normalized reference temperature values in $[0,1]$ for $\lambda_{R}=1$.
A.3 SEJONG Dataset
The SEJONG dataset illustrates the impact of various disguises, including glasses, wigs, and fake beards. Each subject in the dataset is presented with different disguises. As a result, it includes a lot of clothing and hairstyle variations. Most participants have a Southeast Asian or Central Asian ethnic background, but people of other ethnicities are included too. The number of participants identifying as male or female is balanced. This makes it an attractive dataset candidate for thermalization training that is supposed to generalize to a large variety of subjects. Due to reasons of data privacy, we refer to the original publication [7] and abstain from showing SEJONG images.
A.4 Extended Thermalization Ablation Study
Figure 5: T-Fake samples with original images (first row), ‘cold’ images’ (second row), and ‘warm’ images’ (third row). Note the diminished contrasts of the noses and the checks.
Given that the regularizers are solely defined for the synthetic images, we fix $\lambda_{T}=1$ to ensure that the regularizer is on the same scale as the MSE of the real images and vary $\lambda_{W}$ for our ablation. We train all models with the same setup and random seeds. The FID implementation is the default PyTorch-Ignite [15] implementation. Here, we extend the ablation study table in the main document which only displays the best result for $\lambda_{W}=0.01C$. In particular, we display the results of a grid search for $\lambda_{W}$ and $\lambda_{R}$ in Table 7. As described in the main document, we calculate the mean FID and its standard deviation for twenty different subsets of the T-FAKE dataset. Moreover, we use the same setup to compare the perceptual quality of the ‘cold’ and the ‘warm’ setup of our final T-FAKE dataset, see Table 8. Figure 5 shows T-FAKE samples with the ‘cold’ and ‘warm’ variants. In addition, we present some samples generated with different regularization configurations in Figure 6 for the ‘cold’ setup and in Figure 7 for the ‘warm’ setup. Here, we can visually see the impact of the different regularizers.
Regularization $\lambda_{R}=1$ $\lambda_{R}=0$
$\lambda_{W}=1C$ 0.2714 $\pm$ 0.0045 0.3139 $\pm$ 0.0035
$\lambda_{W}=0.1C$ 0.2710 $\pm$ 0.0042 0.3134 $\pm$ 0.0048
$\lambda_{W}=0.01C$ 0.2666 $\pm$ 0.0037 0.3120 $\pm$ 0.0043
$\lambda_{W}=0$ 0.2719 $\pm$ 0.0054 0.3325 $\pm$ 0.0039
Table 7: Impact of the regularization parameters on the perceptual quality measured with the FID with $C=(5\cdot 8^{2})^{-1}$.
According to the FID, both regularizers have a positive impact. The segmentation-based regularizer greatly boosts the perceptual quality, while the effect of the patch-based regularizer is smaller. The optimal FID value is obtained for $\lambda_{W}=0.01C$ and $\lambda_{T}=1$, the parameters of our final model. The perceptual quality of the figures seems in line with the FID. A closer look at the last two rows in both figures shows that the segmentation-based regularizer alone can lead to smoothed facial areas and overly exaggerated differences on the edges of the semantic segmentation. This becomes more apparent for the ‘cold’ setup as it leads to more thermal contrast within the face, see Figure 6. The FID shows only a small difference between the ‘cold’ and ‘warm’ variants. The ‘warm’ variant displays slightly lower FID values. For a visual comparison of the T-FAKE images with real thermal images, we refer to Fig. 8.
Setup ‘Cold’ ‘Warm’
FID 0.2637 $\pm$ 0.0044 0.2699 $\pm$ 0.0078
Table 8: Perceptual comparison of thermal setups ‘cold’ and ‘warm’ using the FID.
Figure 6: Regularization impact on images for ‘cold’ setup: No regularization ($\lambda_{W}=0$, $\lambda_{R}=0$), only patch-based ($\lambda_{W}=0.001C$, $\lambda_{R}=0$), only segmentation-based ($\lambda_{W}=0$, $\lambda_{R}=1$), and final model ($\lambda_{W}=0.001C$, $\lambda_{R}=1$) (top to bottom).
Figure 7: Regularization impact on images for ‘warm’ setup: No regularization ($\lambda_{W}=0$, $\lambda_{R}=0$), only patch-based ($\lambda_{W}=0.001C$, $\lambda_{R}=0$), only segmentation-based ($\lambda_{W}=0$, $\lambda_{R}=1$), and final model ($\lambda_{W}=0.001C$, $\lambda_{R}=1$) (top to bottom).
Appendix B Landmarking
B.1 CHARLOTTE Dataset
The CHARLOTTE dataset contains thermal images with varying thermal conditions, various head positions, and multiple camera distances. Moreover, it contains information about the thermal sensation of the subjects. We refer to Fig. 8 for a visualization of some thermal CHARLOTTE images without landmarks.
Figure 8: CHARLOTTE image samples with different resolutions, environmental conditions, and subjects.
B.2 Landmarking Implementation Details
For training the landmarker $T_{\psi}$, we finetune a model that has been pre-trained on the original FAKE dataset for the sparse landmarker. For the dense landmarker, pre-training happened on the original FAKE dataset including a first version of the T-FAKE dataset as augmentation with a probability of 0.2. During refinement, thermal images are used with a probability of 0.4 split with equal probability for ‘cold’ and ‘warm’ conditions. We finetune for 100 epochs with a learning rate of 0.0004, Adam optimizer with weight decay, batch size of 512, and OneCycleLR scheduler.
B.3 Label Adaptation Implementation Details
Label Adaptation was trained for each method on the predictions on all detected faces on a random 1000-image CHARLOTTE split. We train a model $T_{\zeta}$ for 2000 epochs with a learning rate of 0.002, OneCycleLR, and Adam optimizer. As landmark augmentations, we apply random rotation up to 45° as well as random shearing during training. The label adaptation network is a five-layer perceptron with fully connected layers that takes the predicted landmarks together with the resize factor as input. The latter accounts for varying degrees of quantization at different image sizes in the CHARLOTTE ground truth. The label adaptation generally handles even outlier predictions but can also contain fail cases, (see Figure 10).
RGB Model Inference. We use two different pre-processing approaches to include landmarkers solely developed for RGB images in the evaluation. Firstly, gray-value images, where the temperature between 20° and 45° is normalized and, secondly the pre-processing stack proposed in [12]. The pre-processing stack consists of temperature clamping between 20°C and 45°C, unsharp masking with two sets of parameters with and without temperature inversion. The reported landmarks are the averages over all detected faces. This simple pre-processing stack is a simple method for boosting RGB landmarker performance for thermal images [12]. As a result, we can include RGB landmarkers as a baseline for thermal landmarking models.
B.4 Inference Ablation
We analyze the impact of our inference strategy, see Table 9. We compare inference on i.) the complete image scaled to $224\times 224$ (whole image), ii.) followed by refinement on a bounding box computed from the predictions obtained with i.) and finally iii.) with the sliding window approach described in the main document. Method iii.) produces the best results over all images except for Charlotte low, where ii.) performs slightly better while at the same time being also suitable for real-time estimation.
Figure 9: Results on examples from the CHARLOTTE dataset [3] with different RGB and thermal predictors and our models. For images without landmarks, no faces were detected. The performance of RGB methods can be greatly improved when the images are inverted or sharpened indicated by *. The first column shows the limitations in the CHARLOTTE ground truth: profile annotation convention for frontal views (1st row), quantization artifacts for low-resolution images (4th row), and translated annotations (last row).
Method Charlotte High Charlotte Low Charlotte Side Charlotte Front Charlotte Full
Sliding window ($\bar{\sigma}<\infty$) 0.0740 0.1346 0.0684 0.1420 0.1051
Refined bounding box ($\bar{\sigma}<\infty$) 0.0847 0.1329 0.0696 0.1494 0.1095
Whole image ($\bar{\sigma}<\infty$) 0.1091 0.1868 0.0842 0.2139 0.1490
Table 9: Different strategies for landmark computation. The final results are estimated with sliding windows similar to [53], however, we achieved comparable results when computing the landmarks on input images rescaled to $224\times 224$. Here, we do not exclude high-uncertainty images and evaluate all images without failure rate, i.e., $\bar{\sigma}<\infty$.
Appendix C Large-Scale Visualization
For a large number of T-FAKE samples, we refer to Figures 11 and 12. Here, we simply use the first 128 images based on the numerical naming convention of the original FAKE dataset.
Appendix D Thermal Semantic Segmentation Dataset
Note that by design detailed segmentation masks are available for all T-FAKE images. While the training of a semantic segmentation model was out-of-scope for our work, we want to highlight the fact that our dataset can also be used for such training. The possibility of such segmentation training with synthetic data has already been demonstrated by Wood et al. [52].
Appendix E Limitations
Thermalization. This work depends on the thermalization of the final renders in the FAKE dataset. The dataset contains very difficult lighting conditions and scene compositions that make it powerful to train landmarkers but also made the thermalization particularly challenging and could only be solved with advanced regularization approaches. Despite a good perceptual result, some T-FAKE images contain artifacts, e.g., black patches on the upper body. Nevertheless, these artifacts remain limited.
Dense Landmarks. In this work, we relied on the ground truth 70-point landmarks of the FAKE dataset [52] and dense Mediapipe [35] annotations. Training with the original 320-point and 702-point landmarks could further boost accuracy and lead to an even denser landmarker. However, these landmarks are not publicly available. Furthermore, we only evaluate a mobilenet backbone. Performance for difficult poses and face variability could be achieved with denser models [53].
CHARLOTTE. The CHARLOTTE [3] dataset is among the largest datasets with thermal recordings of faces that contain different levels of image quality as well as has a high variability in poses such as side profile pictures and tilting which makes the dataset ideal as a benchmark. However, the annotation uses a non-standard convention where side profile images have a different number of landmarks than frontal faces. Furthermore, landmarks of low-resolution images are quantized and there are examples of shifted ground truth annotations (see Figure 9).
Figure 10: Label adaptation examples. For side profile views in the CHARLOTTE dataset only landmarks for visible parts of the face exist and additional positions on the forehead are marked. Label adaptation translates the three different landmark conventions we use for evaluation (bottom) into the CHARLOTTE convention (top).
Label Adaptation. We retrain the label adaptation networks for each tested landmarker on its original predictions. Hence, for a given method, a high failure rate on CHARLOTTE means that fewer training images are available for that method. Also, very poor predictions on some of the images such as profile images without failure produce a low quality of adapted landmarks (see Figure 10, TL54 dlib, column 3). It is important to note that the label adaptation does not use visual information from the benchmark dataset to translate landmarks. However, the same individuals were present in both the test and the training dataset splits which might allow the network to learn the facial statistics of individuals. This might give an advantage to landmarkers with only a few landmarks, e.g., TFW [32] (see Figure 10, TFW). The high 68-point landmark performance of TFW on the CHARLOTTE thermal data does not necessarily mean that this generalizes and the advantage of increasing the number of predicted landmarks has been demonstrated for RGB images [53].
Appendix F Societal Impact Statement
An increasing number of risks are associated with collecting and using biometric information, such as (thermal) facial images. As for collecting such information in general, biometric datasets contain information that could be misused. Biometric applications, such as facial recognition or landmarking technology, may have internal biases and may display higher error rates for certain populations than for others in terms of, e.g., ethnicity, gender, or age. The use of synthetic data can be very helpful for such challenges as it allows for explicit control of dataset biases and a decreased need for open-source biometric personal information. Nevertheless, facial recognition technologies can be misused for surveillance purposes and criminal activities. We hope that our research is not employed with malicious intent. Therefore, we strongly encourage the legislative regularization of face recognition and face analysis applications.
Nevertheless, we want to point out the advantages of facial analysis in the thermal domain since this domain is nearly invariant to lighting conditions and inherently difficult to falsify. Highly accurate thermal landmarks can offer security and authentication benefits to society. The computation of reliable thermal landmarks is a key technology for medical screening of various diseases and remote assessment of the physiological and psychophysiological state of humans [12]. Contactless screening for infectious diseases has the potential to reduce environmental impact as well as to increase security through upscaled screening in pandemic or endemic scenarios. This, however, requires a detailed understanding of the limitations of such methods and can otherwise result in a false sense of security or discrimination against false positive users and dense landmarks offer a way to better interpret imaged temperatures from the face and pinpoint contributions from distinct facial areas.
The availability of dense thermal imaging to the biomedical research communities promises new applications and new scientific insights from highly accurate temperature analysis of the face. Traditionally, studies involving thermal imaging of faces are analyzed within a semi-automated pipeline which would require manual region-of-interest definitions [21] that limit the application of the results in out-of-lab environments.
Figure 11: The first 128 FAKE (with removed background) and T-FAKE images with a random choice between the ‘cold’ and the ‘warm’ variant.
Figure 12: The sparse landmarks for the T-FAKE images in Fig. 11